# Unexpected fault activation in underground gas storage. Part II: Definition of safe operational bandwidths

### v1

@@ 6.1 Definition of safe operational bandwidths
After the end of primary production, when the natural fluid pressure $P_{i}$ is reduced to $P_{min,PP}$, the reservoir experiences a relatively fast pressure recovery to $P_{max,CGI/ST}$ during cushion gas injection or gas storage (CO2 or N2). Afterward, UGS or UHS are characterized by a cyclic pressure fluctuation between $P_{min,UGS/UHG}$ and $P_{max,UGS/UHS}$ at the end of the production and injection phases, respectively. In the usual practice and in alignment with the legislation of some countries, e.g., the Netherlands, $P_{max,CGI/ST}\simeq P_{max,UGS/UHS}\simeq P_{i}$ and $P_{min,UGS/UHS}\geq P_{min,PP}$ (Figure 24a). In this framework, guidelines for the definition of “safe operation bandwidths” in gas storage, i.e., operations with a reduced risk of fault reactivation, must identify proper values for the aforementioned pressure bounds. Nevertheless, it has to be recalled that a fault reactivation could always occur aseismically (Cappa et al., 2019) .
The interpretation of the modeling results allows to outline some key guidelines. They are necessarily qualitative because of the theoretical/general framework of the modeling application and the quasi-static nature of the implemented model, which properly simulates the possible inception of fault slip but not seismicity. The numerous scenarios investigated within the study have clearly revealed that fault failure is more likely to happen during CGI/ST and UGS/UHS in depleted reservoirs when:
@@
    fault reactivation is occurred during PP at pressure $P_{seis,PP}$ (Figure 24);
@@
    the reservoir pressure approaches $P_{max,CGI/ST}$, $P_{max,UGS/UHS}$, or $P_{min,UGS/UHS}$.
The resulting guidelines, which link $P_{max,CGI/ST}$, $P_{max,UGS/UHS}$, and $P_{min,UGS/UHS}$ to $P_{min,PP}$ and $P_{seis,PP}$, can be summarized as follows.
Figure 24: (a) Pressure values to be accounted for in the definition of operational safety guidelines. (b) and (c) Safe pressure variation during CGI/ST and UGS/UHS phases in the case of a seismic event occurring during PP at pressure close to $P_{i}$ and $P_{min,PP}$, respectively.
@@
    If a fault reactivation occurs during PP, the pressure change $\Delta P$ spanned during CGI/ST and UGS/UHS phases should be kept smaller than $|P_{i}-P_{min,PP}|$. Indeed, a number of investigated scenarios (see for example Figure 9) reveals that fault activation during primary production leads to a stress redistribution and a new deformed “equilibrated” configuration that is newly loaded, in the opposite direction, when the pressure variation changes the sign. A reasonable rule is to keep $\Delta P$ smaller than maximum between $|P_{seis,PP}-P_{min,PP}|$ (case a) and $|P_{i}-P_{seis,PP}|$ (case b). Generally, case a) or case b) prevails depending on the value of $P_{seis,PP}$ with respect to $P_{i}$ and $P_{min,PP}$. (Figure 24b,c).
@@
    In case a, $P_{max,CGI/ST}$ and $P_{max,UGS/UHS}$ should be kept below $P_{i}$. A reasonable rule of thumb could be to keep $P_{max,CGI/ST}$ and $P_{max,UGS/UHS}$ smaller than the $P_{seis,PP}$.
@@
    In case b, $P_{min,UGS/UHS}$ should be kept above $P_{min,PP}$. A reasonable rule of thumb could be to keep $P_{min,UGS/UHS}$ larger than the $P_{seis,PP}$.
@@
    If during PP activation occurs on a fault separating two reservoir compartments, during CGI/ST or UGS/UHS the pressure difference between the adjacent blocks should be safely kept smaller than $P_{i}-P_{seis,PP}$.
@@
    If no activation occurs during PP, $P_{max,CGI/ST}$ can equate $P_{i}$ with no particular risk of unexpected events during CGI and UGS. Moreover, $\Delta P$ during UGS/UHS can safely span the whole pressure change between $P_{i}$ and $P_{min,PP}$. In fact, the system lies in reloading conditions and behaves practically in an elastic way within the pressure range already experienced.
@@
It has to be emphasized that fault reactivation is largely case-sensitive, depending on the geometry of the fault/reservoir (e.g., the presence of sloped faults, dislocation of the reservoir compartments), differential pore pressure between adjacent reservoir compartments and within each reservoir block, the geomechanical properties of the faults and the reservoir, caprock, under- and overburden. Therefore, specific investigations are fundamental to characterize the actual reservoir setting and quantify more specific bounds. Moreover, the need of more specific analyses holds in relation to the geochemical effects on fault mechanical properties caused by the substitution of formation fluids with CO2, H2 or N2. In the modeling approach used in this work, these possible effects on faults have been neglected due to the lack of laboratory testing. Finally, we underline that the above rules are aimed at avoiding, or at least limiting the probability, of fault activation. For example, according to Wassing et al. (2017), “after the onset of fault reactivation, a further pore pressure decline of 3.7 MPa (or 1.6 MPa with no fault offset) and a critically stressed length of approximately 30 m is needed for the nucleation of a seismic event”. Therefore, since fault reactivation can develop with aseismic slip, these can be regarded as conservative recommendations.
7 Conclusions
@@
Depleted gas reservoirs serve not only as storage facilities for natural gas but also for CO2, N2, and H2. The injection and extraction of various fluids induce changes in the hydraulic and mechanical state of the reservoir at depth. Consequently, deformations of various magnitude occur beneath the subsurface, potentially leading to induced seismic events along fault planes. While the majority of human-induced seismic activities can be explained by fluid injection at pressures beyond the initial levels, triggering the shear stress along fracture surfaces to reach the limit strength, certain recorded events do not fall into this explanation. These particular seismic events, usually called “unexpected”, occur when the pressure falls within the range experienced during primary production.
The main aim of this work is to understand whether and under which circumstances faults can be reactivated (and seismic events possibly induced) during UGS activities. A one-way coupled strategy is adopted to deal with the fluid-structure interaction. The study has been carried out using the typical geological setting of UGS reservoirs in the Netherlands and four underground storage systems, i.e., cyclic storage of CH4 and H2 and permanent storage of CO2 and N2, have been analyzed. This objective is achieved by investigating which are the main factors that control the reactivation of faults and seismicity under permanent or cyclic storage conditions of the stored gases, and to which extent the mechanisms for seismicity are similar to the mechanisms during natural gas storage. Additionally, the study evaluates the potential risk posed by geochemical effects.
The effects of CO2 injection on mechanical properties of faults are negligible as both weakening and hardening effects have been simulated on reservoir and caprock with a ±$\pm 30\%$ change in stiffness moduli with no significant change in fault stability. Instead, the fault critical conditions are controlled by the extreme values of pore-pressure (minimum and maximum) in the formation. Since these values are similar for the natural gas storage operations, the mechanisms for reactivation (which can be aseismic) are also similar. The critical factors (drivers/settings) influencing fault activation during CGI and UGS cycles arrange differently for the boundary and central faults. On faults bounding the reservoir, during UGS the stability is mainly jeopardized by an initial stress regime with small horizontal principal components, failure criterion with low friction angle, large pressure change because of production/injection, and significant contrast between the reservoir and the over- side- and underburden stiffness. If the central fault is on vertical position and allows for symmetry conditions, the major influencing drivers to fault instability are given by geometrical parameters characterizing the fault/reservoir system. Offset of the reservoir compartments, non-vertical fault plane, different pressure change in the two compartments, and fault dip are the features threatening the stability of this fault. Finally, it is worth to highlight the lack of quantitative data regarding N2 and H2 effects on rock properties, and thus on fault stability.
These considerations are supported by the ranking of the combined scenarios. A re-assessment of the project’s findings on geochemical effects causing mechanical weakening should be made at the point where more data from laboratory experiments and case studies becomes available.
CRediT authorship contribution statement
Selena Baldan: Investigation, Formal analysis, Writing - original draft. Massimiliano Ferronato: Conceptualization, Methodology, Writing - review and editing, Supervision. Andrea Franceschini: Methodology, Software, Writing - review and editing. Carlo Janna: Software. Claudia Zoccarato: Conceptualization, Methodology, Formal analysis, Investigation. Matteo Frigo: Software, Investigation. Giovanni Isotton: Conceptualization, Methodology, Formal analysis, Software, Funding acquisition. Cristiano Collettini: Formal analysis, Writing - review and editing. Chiara Deangeli: Methodology, Formal analysis, Writing - review and editing. Vera Rocca: Methodology, Formal analysis, Writing - review and editing. Francesca Verga: Methodology, Formal analysis, Writing - review and editing. Pietro Teatini: Conceptualization, Methodology, Formal analysis, Writing - review and editing, Supervision, Funding acquisition.
@@ Acknowledgements
This research was supported by the State Supervision of Mines (SodM), Ministry of Economic Affairs (The Netherlands), projects KEM01 “Safe Operational Bandwidth of Gas Storage Reservoirs” and KEM39 “Study within the Mining Effects Knowledge Program (KEM-39) on the cyclic storage of gases in the Netherlands”. M.F., A.F., and C.J. from the University of Padova are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). Computational resources were provided by University of Padova Strategic Research Infrastructure Grant 2017: “CAPRI: Calcolo ad Alte Prestazioni per la Ricerca e l’Innovazione”.

### v4

@@ 4.2 Definition of safe operational bandwidths
After the end of PP, when the natural fluid pressure $P_{i}$ is reduced to $P_{min,PP}$, the reservoir experiences a relatively fast pressure recovery to $P_{max,CGI/CCS}$ during CGI or gas storage (CO2 or N2). Afterward, UGS or UHS are characterized by a cyclic pressure fluctuation between $P_{min,UGS/UHG}$ and $P_{max,UGS/UHS}$ at the end of the production and injection phases, respectively. In the usual practice and in alignment with the legislation of some countries, e.g., the Netherlands (Ministerie van Economische Zaken en Klimaat, 2022; TNO, 2018), $P_{max,CGI/CCS}\simeq P_{max,UGS/UHS}\simeq P_{i}$ and $P_{min,UGS/UHS}\geq P_{min,PP}$ (Figure 22-left). In this framework, guidelines for the definition of “safe operation bandwidths” in gas storage, i.e., operations with a reduced risk of fault reactivation, must identify proper values for the aforementioned pressure bounds. Nevertheless, it has to be recalled that a fault reactivation could always occur aseismically (Cappa et al., 2019).
The interpretation of the modeling results allows outlining some key guidelines. When a reservoir is depleted during PP, the faults adjust to a new, permanently deformed configuration. If the same faults slipped during PP, they become more sensitive to later pressure changes, because the stress state no longer returns to its original configuration when pressure is increased again. This means that even pressure variations that remain within the range already experienced during PP can bring a fault back close to reactivation. The safe operating bandwidth therefore depends not only on the absolute pressure values, but also on how much the pressure changes with respect to the amounts that previously caused (or did not cause) slip. A physically consistent operational rule is therefore to limit the pressure excursions during CGI/CCS and UGS/UHS so that they do not exceed the pressure changes that were sustained during PP without inducing slip. The following operational guidelines summarize how this principle translates into practical limits for pressure cycling in depleted gas reservoirs.
They are necessarily qualitative because of the theoretical/general framework of the modeling application and the quasi-static nature of the implemented model, which properly simulates the possible inception of fault slip but not the seismic evolution. The numerous scenarios investigated within the study have clearly revealed that fault failure is more likely to happen during CGI/CCS and UGS/UHS in depleted reservoirs when:
@@
    fault reactivation is occurred during PP. We refer to the pressure of seismic occurrence as $P_{seis,PP}$. As an example, in scenario #4b faults are not active during PP and remain far from critical conditions during CGI and UGS as well (Figure 11 and Table 5);
@@
    the reservoir pressure approaches $P_{max,CGI/ST}$, $P_{max,UGS/UHS}$, or $P_{min,UGS/UHS}$ (e.g., Figures 17 and 20).
The resulting guidelines, which link $P_{max,CGI/CCS}$, $P_{max,UGS/UHS}$, and $P_{min,UGS/UHS}$ to $P_{min,PP}$ and $P_{seis,PP}$, can be stated as follows. The outcome of scenario C2 is particularly illustrative in this regard (Figures 17 and 20).
Figure 22: Left: pressure values to be accounted for in the definition of operational safety guidelines. Middle and right: safe pressure variation during CGI, CCS, and UGS/UHS phases in the case of a seismic event occurring during PP at pressure close to $P_{i}$ and $P_{min,PP}$, respectively.
@@
    If a fault reactivation occurs during PP, the pressure change $\Delta P$ spanned during CGI/CCS and UGS/UHS phases should be kept smaller than $|P_{i}-P_{min,PP}|$. Indeed, a number of investigated scenarios reveals that fault activation during PP leads to a stress redistribution and a new deformed “balanced” configuration that is newly loaded, in the opposite direction, when the pressure variation changes its sign. This behavior is clearly observed in Scenario #2 (Figure 8), Scenario C2 (Figures 17 and  20), and also in Scenarios #3 and #4a, where the fault approaches criticality during CGI despite $p<P_{i}$. A reasonable rule is to keep $\Delta P$ smaller than maximum between $|P_{seis,PP}-P_{min,PP}|$ (range highlighted with a in the middle panel of Figure 22) and $|P_{i}-P_{seis,PP}|$ (range highlighted with b in the right panel of Figure 22), i.e., the maximum pressure difference experienced by the reservoir during PP without the occurrence of a seisimic event. The classification into these two classes is determined by the value of $P_{seis,PP}$ in relation to $P_{i}$ and $P_{min,PP}$, as illustrated in Figure 22, particularly in the middle and right panels.
@@
    In case a, represented in the middle panel of Figure 22, $P_{max,CGI/CCS}$ and $P_{max,UGS/UHS}$ should be kept below $P_{i}$. A reasonable rule of thumb could be to keep $P_{max,CGI/CCS}$ and $P_{max,UGS/UHS}$ smaller than $P_{seis,PP}$.
@@
    In case b, right panel of Figure 22, $P_{min,UGS/UHS}$ should be kept above $P_{min,PP}$. A reasonable rule of thumb could be to keep $P_{min,UGS/UHS}$ larger than $P_{seis,PP}$.
@@
    If during PP activation occurs on a fault that separates two reservoir compartments, during CGI/CCS or UGS/UHS the pressure difference between adjacent blocks should be kept safely less than $P_{i}-P_{seis,PP}$ (scenario #5b, Figure 14).
@@
    If no activation occurs during PP, $P_{max,CGI/CCS}$ can equate $P_{i}$ with no particular risk of unexpected events during CGI and UGS. Moreover, $\Delta P$ during UGS/UHS can safely span the whole pressure change between $P_{i}$ and $P_{min,PP}$. In fact, the system operates under reloading conditions and exhibits predominantly elastic behavior within the pressure range experienced previously (scenario #4b, Figure 11).
A graphical representation of the guidelines is provided in Figure 23.
Figure 23: Graphycal representation of the operational guidelines for a compartmentalized, previously produced gas reservoir. Aa key factor is represented by the pressure at which fault reactivation occurred during PP.
4.3 Limitations of the current modeling approach
The modeling framework adopted in this study involves several simplifications that are necessary to address the physical problem at the scale of real-world gas reservoirs. The principal limitations are discussed in this section.
Although the model setup is based on representative geological configurations, both the fault-reservoir geometry and the spatial distribution of physical properties are simplified to allow the problem to be computationally manageable at the basin scale. Geometries are schematized and material properties are homogeneously assigned per domain block. As a result, the model does not capture fine-scale heterogeneities, such as local discontinuities, grain-scale damage zones, or small-scale structural complexity. It is crucial to highlight that fault reactivation is largely site-sensitive, depending on the geometry of the fault/reservoir (e.g., the presence of sloped faults, dislocation of the reservoir compartments), differential pore pressure between adjacent reservoir compartments and within each reservoir block, the geomechanical properties of the faults and the reservoir, caprock, under- and over-burden. Therefore, specific investigations are of paramount importance to characterize the actual reservoir setting and quantify more specific bounds.
For CH4 and H2, the loading history follows a synthetic temporal evolution, designed to replicate typical seasonal injection–withdrawal amplitudes observed in Dutch UGS operations. Unlike actual field conditions, where pressure amplitudes tend to increase gradually over time, our model imposes high-amplitude cycles from the beginning of UGS/UHS ($\sim$50% of the depletion range) (Figure 2, bottom-right panel). This idealization is consistent with long-term UGS behavior (see, e.g., TNO (2015)) and allows us to evaluate the fault response under representative upper-bound stress conditions.
The simulations adopt a one-way hydro-mechanical coupling strategy, where the pore pressure $p$ is provided by a flow model and its contribution for the geomechanical model is regarded as an external source of strength. This modeling choice is recognized as an approximation: it may overestimate pressure buildup near faults and anticipate fault reactivation compared to fully coupled simulations, especially under long-term or cyclic loading. Previous studies have shown that the divergence between one-way and two-way coupling can increase significantly over time (Prevost, 2013). However, the use of one-way coupling in this study is a practical compromise and it is fully warranted on the space and time scale of interest (Buijze et al., 2019; Gambolati et al., 2000). It enables the execution of a large number of 3D, elasto-plastic simulations with a fault system and realistic boundary conditions, while still capturing the relative sensitivity of fault stability to pressure changes.
The model adopts a quasi-static formulation and does not include inertial effects or dynamic rupture propagation (Franceschini et al., 2024). As a result, it cannot distinguish between aseismic and unstable seismic fault slip, nor simulate the nucleation and growth of dynamic instabilities. Fault activation is evaluated based on a static failure criterion (in our case, Coulomb’s formulation), and is used here solely as an indicator of the onset of slip. The model is therefore intended to identify critical loading conditions for reactivation, not to reproduce the full physics of earthquake ruptures.
Building on the previous point, the operational guidelines derived from our simulations are formulated as conservative thresholds that aim to reduce the likelihood of any fault reactivation, regardless of whether slip would manifest aseismically or seismically. It is worth noting that these guidelines are not universally applicable: they are restricted to a specific class of reservoirs, namely previously depleted gas fields with well-defined geometries and pressure histories, as commonly found in the Dutch subsurface. Extrapolation to other types of storage sites would require additional modeling and site-specific calibration.
@@ 4.3 Limitations of the current modeling approach
Moreover, the need of more specific analyses holds in relation to the geochemical effects on fault mechanical properties caused by the substitution of formation fluids with CO2, H2, or N2. In the modeling approach used in this work, these potential effects on faults have only been superficially addressed, due to the lack of laboratory testing. Finally, we underline that the above rules are aimed at avoiding, or at least limiting, the probability of fault activation. For example, according to Wassing et al. (2017), “after the onset of fault reactivation, a further pore pressure decline of 3.7 MPa (or 1.6 MPa with no fault offset) and a critically stressed length of approximately 30 m is needed for the nucleation of a seismic event”. Therefore, since fault reactivation can develop with aseismic slip, the recommendations listed above can be considered conservative.
Another important limitation concerns the treatment of pore pressure in the seal and underburden. In the current modeling framework, these units are not included in the flow simulation and are assumed to remain at their initial hydrostatic pressure throughout all production and storage stages. However, pressure diffusion into the seal and underburden would act to smooth pressure gradients at the reservoir boundaries, thereby reducing stress localization and mitigating the development of stress singularities at fault–layer intersections. Moreover, spatial pressure heterogeneities and locally elevated injection pressures that may develop near wells during CGI and UGS are not represented. Thermo-mechanical effects associated with cooling during gas injection are also neglected.
In the present modeling framework, reservoir deformation is represented using a simplified mechanical description that does not explicitly account for elasto-plastic or visco-plastic deformation, strain hardening, or time-dependent effects. Laboratory studies on Slochteren sandstone show that inelastic deformation contributes significantly to reservoir compaction already at small strains and leads to stress-path hysteresis, with different stress evolution during loading and unloading phases (Pijnenburg et al., 2019b). As a consequence, the stress paths modeled here during pressure depletion and subsequent pressurization are largely elastic and reversible, and do not capture the irreversible stress-path shifts documented experimentally. Including inelastic and time-dependent reservoir behavior would likely modify the magnitude and timing of stress transfer to faults, particularly by reducing elastic stress recovery during pressurization and by shifting the conditions at which fault criticality is reached.
4.4 Conclusions and future prospects
@@
Depleted gas reservoirs can serve not only as storage facilities for natural gas (i.e., CH4) but also for CO2, N2, and H2. The injection and/or withdrawal of these fluids induce changes in the hydraulic and mechanical state of the reservoir. Consequently, deformations and variations in the stress regime develop beneath the subsurface, potentially leading to (seismic/aseismic) reactivation of faults near the reservoir. While the majority of human-induced seismic activities can be associated to injection or withdrawal of fluids at pressures above the original formation pressure causing significant pressure decline (Ellsworth, 2013), which trigger shear stress along faults to reach their limit strength, a few recorded events do not fit this explanation. These seismic events, somehow “unexpected”, develop in reservoirs where:
  • 1.
    seismic events already occurred during the primary production phase;
  • 2.
    in correspondence to a pressure value already experienced by the reservoir during PP.
The simulation of various realistic scenarios have allowed to define the critical factors influencing fault activation during CGI/CCS and UGS/UHS cycles. The stability of the faults bounding the reservoir compartments is mainly jeopardized by an initial stress regime with small horizontal principal components, low friction angle, large pressure change because of injection/withdrawal, and significant contrast between the reservoir and the over-, side-, and underburden stiffness. Concerning faults located between producing blocks, the drivers mainly influencing fault instability are the geometrical setting of the fault/reservoir system, i.e., the offset between reservoir compartments and the fault dip angle, together with the different pressure change in adjacent compartments.
Notice that the pressure recovery and drop addressed in the simulations have been defined based on regulations in the Netherlands, which do not allow the pressure to rise above the initial natural value. Due to this constrain, the mechanisms causing fault reactivation remain similar, irrespective of the fluid considered. The possible mechanical weakening and hardening of the reservoir associated to non-natural pore fluids interaction with the solid grains does not impact significantly the outcomes.
The modeling outcomes have enabled the formulation of general guidelines to define safe operational bandwidths for USS sites, specifically the pressure range over which “unexpected” seismic event can be excluded. These minimum and maximum pressure thresholds are closely linked to the pressures at which seismic events occurred during primary production. The occurrence of a seismic event during PP provides valuable insights for delineating pressure bandwidths within which fault reactivation is highly unlikely during UGS, UHS, CGI, and CCS activities.
These conservative recommendations can serve as a preliminary methodology for reducing the potential risks associated to seismic activity in similar contexts. However, more specific and in-depth evaluations must follow, taking into account the peculiarities of each individual case study.
5  Supplementary Materials
Figure 24: Scenario #1 ("Reference"): depth profiles of the criticality index $\chi$ on all faults at the end of PP (loading step 10). High criticality develops at the reservoir top and bottom.
@@ Acknowledgements
This research was supported by the State Supervision of Mines (SodM), Ministry of Economic Affairs (The Netherlands), projects KEM01 “Safe Operational Bandwidth of Gas Storage Reservoirs” and KEM39 “Study within the Mining Effects Knowledge Program (KEM-39) on the cyclic storage of gases in the Netherlands”. M.F., A.F., and C.J. from the University of Padova are members of the Gruppo Nazionale Calcolo Scientifico – Istituto Nazionale di Alta Matematica (GNCS-INdAM). Computational resources were provided by University of Padova Strategic Research Infrastructure Grant 2017: “CAPRI: Calcolo ad Alte Prestazioni per la Ricerca e l’Innovazione”.